OpenStudy (anonymous):

Simplify 2 to the 5th over 3 squared all raised to the 4th power. 2 to the 20th over 3 to the 8th 2 to the 9th over 3 to the 8th 8 to the 5th over 12 to the 2nd 2 over 3 to the 2nd

OpenStudy (anonymous):

Please help!!! ;)

OpenStudy (anonymous):

@freckles @phi @perl

OpenStudy (studygurl14):

\(\Large(\frac{2^5}{3^2})^4\) First, find 2^5 and 2^2

OpenStudy (studygurl14):

3^2 I mean

OpenStudy (anonymous):

3 x 2?

OpenStudy (anonymous):

@StudyGurl14 pls help is it multiplying??

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@perl @triciaal @ganeshie8 @Hero @phi

OpenStudy (anonymous):

2 times 5 = 10 and 3 times 2 = 6

OpenStudy (tkhunny):

\(\left(a^{b}\right)^c = a^{b*c}\) Do that.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

sorry but i dont know how @tkhunny

OpenStudy (anonymous):

@StudyGurl14 pls explain

OpenStudy (studygurl14):

3^2 = 3 x 3

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

9

OpenStudy (studygurl14):

yeah.

OpenStudy (tkhunny):

Follow the pattern. \(\left(\dfrac{2^{5}}{3^{2}}\right)^{4} = \dfrac{\left(2^{5}\right)^{4}}{\left(3^{2}\right)^{4}}\) I'm pretty sure you can do that. Now the other part.

OpenStudy (anonymous):

ok what about the other one? is it 2 x 6?

OpenStudy (anonymous):

studygirl

OpenStudy (anonymous):

ugh tkhunny i dont! please its complicated

OpenStudy (tkhunny):

You should not ever see a '9' during the course of this problem.

OpenStudy (anonymous):

what about 6?

OpenStudy (anonymous):

is it 60/40?

OpenStudy (anonymous):

like do u multiply 2 and 5 then 3 and 2 and then both by 4?

OpenStudy (anonymous):

ok start over. first you multiply what?

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

is the final answer 0.66 to the 12th?

OpenStudy (tkhunny):

Use the rule. \(\left(a^{b}\right)^{c} = a^{b*c}\) -- That's it. No great complication. \(\left(2^{5}\right)^{4} = 2^{5*4} = 2^{20}\) -- No more. No less. Done. You do the denominator.